OpenStudy (anonymous):

2 3/8 + 3 3/7 Improper forms = 19/8 + 24/7

OpenStudy (anonymous):

do u wanna see if its right?

OpenStudy (anonymous):

I'd like to solve it. I just made it easier for whoever would like to help. :)

OpenStudy (anonymous):

CAN SOMEONE PLEASE PLEASE PLEASE HELP ME WITH THIS? Two functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = −x + 6 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x).

OpenStudy (anonymous):

oh ok thats fine ok so what do we need to do now

OpenStudy (anonymous):

alexa ill look at ur probelm later

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

k now lelenney any ideas on what we need to do now?

OpenStudy (anonymous):

when adding fractions we need to have common denominators any ideas on whats the common denominator between 7 and 8?

OpenStudy (anonymous):

it would be 56 u know what to do next?

OpenStudy (anonymous):

Sorry, OS was failing to load.

OpenStudy (anonymous):

It took it about five minutes to load.

OpenStudy (anonymous):

24 * 8 19 * 7

OpenStudy (anonymous):

for the first fraction you need to multiply what by 8 to get 56? and after that multiply thr top by the same number as the bottom

OpenStudy (anonymous):

and do that for both fractions

OpenStudy (anonymous):

192 133 56 56

OpenStudy (anonymous):

7. To answer your question.

OpenStudy (anonymous):

So do I just add the numerators now?

OpenStudy (anonymous):

If so, it is now 325/56

OpenStudy (anonymous):

yup because they have a common denominator

OpenStudy (anonymous):

and can u simplify

OpenStudy (anonymous):

May I have assistance on that?

OpenStudy (anonymous):

I severely dislike fractions... However, they must stay fractions..

OpenStudy (anonymous):

sure is there a common factor between 325 and 56?

OpenStudy (anonymous):

Ermm.. Idk.

OpenStudy (anonymous):

multiple choice problem?

OpenStudy (anonymous):

Yeah. 5 45/6 5 11/14

OpenStudy (anonymous):

5 45/56 *

OpenStudy (anonymous):

Those are the ones with the proper whole number.

OpenStudy (anonymous):

5 does not go into 56

OpenStudy (anonymous):

I think it is 5 45/56.. Correct?

OpenStudy (anonymous):

It is correct.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

alexa now ur turn basically what u do is tat u substitute the values in to get the new ones